OpenStudy (anonymous):

If x =3 and y = -2 , then y^4 - 4x = ? A. -28 B. -20 C. -4 D. 4

OpenStudy (anonymous):

4 so D...

OpenStudy (anonymous):



OpenStudy (anonymous):

-2 * -2 = 4 4 * -2 = -8 -8 * -2 = 16 16 - 4(3) = ? 16 - 12 = 4 D

OpenStudy (anonymous):

its not 4 when you have a neg and potive and you subtracting you keep the sign of the 1st number change the minus to addition and take sign of highest number ** (:

OpenStudy (anonymous):

-2^4 does not equal -16. That is why you a getting -28.

OpenStudy (anonymous):

Krewe your ugly :) And thanks guys i got the answer

OpenStudy (anonymous):

Alright, thanks bud.

OpenStudy (anonymous):

gage (: be nice

OpenStudy (anonymous):

lol kk.